# Open-Set Domain Adaptation Under Background Distribution Shift: Challenges and A Provably Efficient Solution

### v2

@@ A.1 Theoretical analysis
$y_{i}\cdot{\mathbf{w}}^{\top}_{\mathrm{\text{DD}}}{\mathbf{x}}_{i}$ $=\left[2\left(r^{2}_{\mu}-\sigma r_{\mu}t\right)^{-1}\cdot\mu-4\sigma^{-2}kd^{-1}\bar{\xi}_{{\mathcal{T}},1}\right]^{\top}\left[\mu+y_{i}\xi_{i}\right]$
$=2(r^{2}_{\mu}-\sigma r_{\mu}t)^{-1}(r^{2}_{\mu}+y_{i}\mu^{\top}\xi_{i})-4\sigma^{-2}kd^{-1}y_{i}\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},1}$
From eq.˜14 we have that $|y_{i}\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},1}|<\sigma^{2}t\sqrt{\frac{d}{k}}$ w.p. at least $1-2\exp\left[-ct^{2}\right]$ for $t\leq d\sqrt{|U|}\sigma^{2}$. Hence, we can say w.p. at least $1-2\exp\left[-ct^{2}\right]$,
$y_{i}\cdot{\mathbf{w}}^{\top}_{\mathrm{\text{DD}}}{\mathbf{x}}_{i}$ $\geq 2(r^{2}_{\mu}-\sigma r_{\mu}t)^{-1}(r^{2}_{\mu}+y_{i}\mu^{\top}\xi_{i})-4\sigma^{-2}kd^{-1}\cdot\sigma^{2}t\sqrt{\frac{d}{k}}$
$=2(r^{2}_{\mu}-\sigma r_{\mu}t)^{-1}(r^{2}_{\mu}+y_{i}\mu^{\top}\xi_{i})-4t\sqrt{\frac{k}{d}}$
$\geq 2(r^{2}_{\mu}-\sigma r_{\mu}t)^{-1}(r^{2}_{\mu}+y_{i}\mu^{\top}\xi_{i})-1\text{ ... (for $k\leq d/(16t^{2})$ and $t>1$)}$
We also have that $|y_{i}\mu^{\top}\xi_{i}|<t\sigma r_{\mu}$ w.p. at least $1-2\exp(-ct^{2}/2)$. Putting these together
@@
Now for $i\in{\mathcal{I}_{{\mathcal{T}},1}}$, we repeat these steps, but this time use eq.˜15 to bound $|\xi^{\top}_{i}\bar{\xi}_{{\mathcal{T}},1}|\geq\frac{\sigma^{2}d}{2k}$ with probability at least $1-3\exp(-ct^{2})$ for $t\leq\frac{\sqrt{d}}{2(2+\sqrt{|U|-1})}$ and $|\eta^{\top}\bar{\xi}_{{\mathcal{T}},1}|<t\sigma r_{\eta}k^{-1}$ w.p. at least $1-2\exp\left(-ct^{2}/2\right)$. We get again that
$y_{i}\cdot{\mathbf{w}}^{\top}_{\mathrm{\text{DD}}}{\mathbf{x}}_{i}$ $=2(r^{2}_{\mu}-\sigma r_{\mu}t)^{-1}(-y_{i}\mu^{\top}\eta+y_{i}\cdot\mu^{\top}\xi_{i})-4\sigma^{-2}kd^{-1}(-y_{i}\eta^{\top}\bar{\xi}_{{\mathcal{T}},1}+y_{i}\cdot\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},1})$
$=4\sigma^{-2}kd^{-1}(-\eta^{\top}\bar{\xi}_{{\mathcal{T}},1}+\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},1})$
@@
Finally, let us calculate ${\mathbf{w}}_{\mathrm{\text{DD}}}^{\top}(\mu-\eta)$,
${\mathbf{w}}_{\mathrm{\text{DD}}}^{\top}(\mu-\eta)$ $=2(r_{\mu}^{2}-\sigma r_{\mu}t)^{-1}r^{2}_{\mu}+4\sigma^{-2}kd^{-1}\eta^{\top}\bar{\xi}_{{\mathcal{T}},1}$
@@
${\mathbf{w}}_{\mathrm{\text{DD}}}^{\top}(\mu-\eta)>0$
@@
$\|{\mathbf{w}}_{\mathrm{\text{DD}}}\|=\left\|\frac{2}{r^{2}_{\mu}-\sigma r_{\mu}t}\cdot\mu-\frac{4k}{\sigma^{2}d}\bar{\xi}_{{\mathcal{T}},1}\right\|.$
@@
$\|{\mathbf{w}}_{\mathrm{\text{DD}}}\|\leq\frac{2}{r^{2}_{\mu}-\sigma r_{\mu}t_{1}}\cdot\|\mu\|+\frac{4k}{\sigma^{2}d}\cdot\|\bar{\xi}_{{\mathcal{T}},1}\|.$
We know that $\|\mu\|=r_{\mu}$ and $\bar{\xi}_{{\mathcal{T}},1}\sim\mathcal{N}(0,\frac{\sigma^{2}}{k}I_{d})$. Hence, using the concentration bound in 12, we have with probability at least $1-\exp{\left(-\frac{t^{2}}{4}\right)}$ that
@@
$\leq\frac{2}{r_{\mu}-\sigma t}+8\sigma^{-1}\sqrt{\frac{k}{d}}\text{ .... for $t\leq\frac{3}{2}\sqrt{d}$}$
@@
$\|{\mathbf{w}}^{*}\|\geq 2\|\eta(\frac{2|U|-k}{|U|})+\bar{\xi}_{U}\|^{-1}$
$\|{\mathbf{w}}^{*}\|\geq 2(\|\eta\|(\frac{2|U|-k}{|U|})+\|\bar{\xi}_{U}\|)^{-1}$
Similarly, using concentration bound in 12 for $\|\bar{\xi}_{U}\|$, we get with probability at least $1-\exp{\left(-\frac{t^{2}}{4}\right)}$,
@@
$C_{d}\log(\frac{c_{d}}{\delta})\leq\frac{\sqrt{d}}{4\sqrt{N_{{\mathcal{S}}}+N_{{\mathcal{T}}}}},\frac{1}{2\sqrt{N_{{\mathcal{T}}}}}\leq r_{\mu}\leq\min\left(2\sqrt{N_{{\mathcal{T}}}}(1-5\tau),16\sqrt{N_{{\mathcal{T}},1}}\tau\right),r_{\eta}\geq\frac{1}{2\sqrt{N_{{\mathcal{T}},1}}},\frac{3}{r_{\eta}}+\frac{3-5\tau}{r_{\mu}}\leq\frac{3}{8}\sqrt{N_{{\mathcal{T}},1}}-2$
then with probability at least $1-\delta$ over the drawing of $(S_{\mathcal{S}},S_{\mathcal{T}})$ as described in problem, $AU-ROC(\hat{{\mathbf{w}}}_{\mathrm{\text{color}}})\geq$... as defined in first part of eq.˜2 is at least $Q\left(\frac{-\tau\sqrt{d}}{\sqrt{2}\|\hat{{\mathbf{w}}}_{\mathrm{\text{color}}}\|}\right)\geq Q\left(\frac{-\tau\sqrt{d}}{\sqrt{2}(\frac{3}{r_{\eta}}+\frac{3-5\tau}{r_{\mu}}+2)}\right)\geq Q\left(\frac{-\tau\sqrt{d}}{\sqrt{2}(1-\tau)}\right)$.
@@
$\mathcal{L}({\mathbf{w}},\lambda,\nu)=\|{\mathbf{w}}\|^{2}-\lambda_{{\mathcal{S}}}^{\top}(\bm{X}_{{\mathcal{S}}}{\mathbf{w}})+\lambda_{{\mathcal{T}}}^{\top}(\mathbf{1}_{{\mathcal{T}}}+\bm{X}_{{\mathcal{T}}}{\mathbf{w}})-\nu\cdot({\mathbf{w}}^{\top}(\mu-\eta)+\tau)$
@@
$\nabla_{{\mathbf{w}}}\mathcal{L}({\mathbf{w}},\lambda,\nu)=2{\mathbf{w}}-\bm{X}^{\top}_{{\mathcal{S}}}\lambda_{{\mathcal{S}}}+\bm{X}^{\top}_{{\mathcal{T}}}\lambda_{{\mathcal{T}}}-\nu\cdot(\mu-\eta)=0$
$\Leftrightarrow{\mathbf{w}}=\frac{1}{2}\left(\nu\cdot(\mu-\eta)+\bm{X}^{\top}_{{\mathcal{S}}}\lambda_{{\mathcal{S}}}-\bm{X}^{\top}_{{\mathcal{T}}}\lambda_{{\mathcal{T}}}\right)$
@@
$\mathcal{L}(\lambda,\nu)=-\nu\cdot\tau+\lambda_{{\mathcal{T}}}^{\top}\mathbf{1}_{{\mathcal{T}}}-\frac{1}{4}\|\nu\cdot(\mu-\eta)+\bm{Z}^{\top}\lambda\|^{2}$
Consider setting $\lambda=\frac{\nu}{|U|}\mathbf{1}_{S_{{\mathcal{T}}}}$, where $U$ has randomly drawn $|U|$ ($|U|$ is to be determined) examples from $S_{{\mathcal{T}}}$. Also set $\nu=\frac{(1-\tau)}{\|\bar{\xi}_{S_{{\mathcal{T}}}}\|^{2}}$. We have
$\bm{Z}^{\top}\lambda$ $=-\nu\cdot\mu+\nu\bar{\xi}_{S_{{\mathcal{T}},0}}+\nu\cdot\eta+\nu\bar{\xi}_{S_{{\mathcal{T}},1}}$
$=\nu(\eta-\mu)+\nu\bar{\xi}_{S_{{\mathcal{T}}}}$
@@
${\mathbf{w}}_{\mathrm{color}}=\alpha\eta+(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}})\mu-\beta\bar{\xi}_{S_{{\mathcal{T}},0}}$
${\mathbf{w}}_{\mathrm{color}}=2\frac{\eta}{r^{2}_{\eta}}+(\frac{2}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}})\mu-\frac{1}{\sigma^{2}}\cdot\frac{|U|-k}{d}\bar{\xi}_{S_{{\mathcal{T}},0}}$
${\mathbf{w}}_{\mathrm{color}}=2\frac{\eta}{r^{2}_{\eta}}+(\frac{2}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}})\mu-(|U|-k)\bar{\xi}_{S_{{\mathcal{T}},0}}$
@@
${\mathbf{w}}_{\mathrm{color}}^{\top}(\eta-\xi_{i})=\alpha r^{2}_{\eta}+\beta\xi_{i}^{\top}\bar{\xi}_{S_{{\mathcal{T}},0}}>1$
@@
${\mathbf{w}}_{\mathrm{color}}^{\top}(-y_{i}\eta+y_{i}\xi_{i})$ $={\mathbf{w}}_{\mathrm{color}}^{\top}(\eta-\xi_{i})$
$=\alpha r^{2}_{\eta}-\alpha\eta^{\top}\xi_{i}+\beta\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},0}$
We know that $|\eta^{\top}\xi_{i}|<\sigma tr_{\eta}$ w.p. at least $1-2\exp\left(-ct^{2}/2\right)$ and $|\xi_{i}\bar{\xi}_{{\mathcal{T}},0}|\leq\sigma^{2}t\sqrt{\frac{d}{|U|-k}}$ w.p. at least $1-2\exp(-ct^{2})$ for $t\leq d\sqrt{|U|}\sigma^{2}$. Hence, w.p. at least $1-2\exp\left(-ct^{2}/2\right)-2\exp(-ct^{2})$, we can say that,
${\mathbf{w}}_{\mathrm{color}}^{\top}(\eta-\xi_{i})\geq\alpha r^{2}_{\eta}-\alpha\sigma tr_{\eta}-\beta\sigma^{2}t\sqrt{\frac{d}{|U|-k}}$
@@
${\mathbf{w}}_{\mathrm{color}}^{\top}(\mu+\xi_{i})=(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}})r^{2}_{\mu}+\beta\xi_{i}^{\top}\bar{\xi}_{S_{{\mathcal{T}},0}}>1$
@@
${\mathbf{w}}_{\mathrm{color}}^{\top}(-y_{i}\mu+y_{i}\xi_{i})$ $={\mathbf{w}}_{\mathrm{color}}^{\top}(\mu-\xi_{i})$
$=\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)(\mu^{\top}\mu-\mu^{\top}\xi_{i})-\beta(\mu^{\top}\bar{\xi}_{{\mathcal{T}},0}-\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},0})$
We know that $|\mu^{\top}\xi_{i}|\leq\sigma tr_{\mu}$ w.p. at least $1-2\exp\left(-ct^{2}/2\right)$ and $|\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},0}|\geq\frac{\sigma^{2}d}{2(|U|-k)}$ w.p. at least $1-2\exp(-ct^{2})$ for $t\leq\sigma^{2}$ and $t\leq\frac{\sqrt{d}}{2(2+\sqrt{|U|-1})}$. We also know that $|\mu^{\top}\bar{\xi}_{{\mathcal{T}},0}|\leq\frac{\sigma tr_{\mu}}{|U|-k}$ w.p. at least $1-2\exp(-ct^{2}/2)$ Hence, w.p. at least $1-4\exp\left(-ct^{2}/2\right)-2\exp(-ct^{2})$, we can say that,
${\mathbf{w}}_{\mathrm{color}}^{\top}(-y_{i}\mu+y_{i}\xi_{i})$ $\geq\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)(r^{2}_{\mu}-\sigma tr_{\mu})+\beta\left(-\frac{\sigma tr_{\mu}}{|U|-k}+\frac{\sigma^{2}d}{2(|U|-k)}\right)$
$\geq\frac{r^{2}_{\mu}}{2}\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)+\beta\left(\frac{\sigma^{2}d-r^{2}_{\mu}}{2(|U|-k)}\right)\text{ ..... for $r_{\mu}\geq 2\sigma t$}$

### v3

@@ A.1 Theoretical analysis
$y_{i}\cdot{\mathbf{w}}^{\top}_{\mathrm{\text{DD}}}{\mathbf{x}}_{i}$ $=\left[2\left(r^{2}_{\mu}-\sigma r_{\mu}t\right)^{-1}\cdot\bm{\mu}-4\sigma^{-2}kd^{-1}\bar{\xi}_{{\mathcal{T}},1}\right]^{\top}\left[\bm{\mu}+y_{i}\xi_{i}\right]$
$=2(r^{2}_{\mu}-\sigma r_{\mu}t)^{-1}(r^{2}_{\mu}+y_{i}\bm{\mu}^{\top}\xi_{i})-4\sigma^{-2}kd^{-1}y_{i}\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},1}$
From eq.˜24 we have that $|y_{i}\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},1}|<\sigma^{2}t\sqrt{\frac{d}{k}}$ w.p. at least $1-2\exp\left[-ct^{2}\right]$ for $t\leq d\sqrt{|U|}\sigma^{2}$. Hence, we can say w.p. at least $1-2\exp\left[-ct^{2}\right]$,
$y_{i}\cdot{\mathbf{w}}^{\top}_{\mathrm{\text{DD}}}{\mathbf{x}}_{i}$ $\geq 2(r^{2}_{\mu}-\sigma r_{\mu}t)^{-1}(r^{2}_{\mu}+y_{i}\bm{\mu}^{\top}\xi_{i})-4\sigma^{-2}kd^{-1}\cdot\sigma^{2}t\sqrt{\frac{d}{k}}$
$=2(r^{2}_{\mu}-\sigma r_{\mu}t)^{-1}(r^{2}_{\mu}+y_{i}\bm{\mu}^{\top}\xi_{i})-4t\sqrt{\frac{k}{d}}$
$\geq 2(r^{2}_{\mu}-\sigma r_{\mu}t)^{-1}(r^{2}_{\mu}+y_{i}\bm{\mu}^{\top}\xi_{i})-1\text{ \hskip 11.38109pt for $k\leq d/(16t^{2})$ and $t>1$}$
We also have that $|y_{i}\bm{\mu}^{\top}\xi_{i}|<t\sigma r_{\mu}$ w.p. at least $1-2\exp(-ct^{2}/2)$. Putting these together
@@
Now for $i\in{\mathcal{I}_{{\mathcal{T}},1}}$, we repeat these steps, but this time use eq.˜25 to bound $|\xi^{\top}_{i}\bar{\xi}_{{\mathcal{T}},1}|\geq\frac{\sigma^{2}d}{2k}$ with probability at least $1-3\exp(-ct^{2})$ for $t\leq\frac{\sqrt{d}}{2(2+\sqrt{|U|-1})}$ and $|\bm{\eta}^{\top}\bar{\xi}_{{\mathcal{T}},1}|<t\sigma r_{\eta}k^{-1}$ w.p. at least $1-2\exp\left(-ct^{2}/2\right)$. We get again that
$y_{i}\cdot{\mathbf{w}}^{\top}_{\mathrm{\text{DD}}}{\mathbf{x}}_{i}$ $=2(r^{2}_{\mu}-\sigma r_{\mu}t)^{-1}(-y_{i}\bm{\mu}^{\top}\bm{\eta}+y_{i}\cdot\bm{\mu}^{\top}\xi_{i})-4\sigma^{-2}kd^{-1}(-y_{i}\bm{\eta}^{\top}\bar{\xi}_{{\mathcal{T}},1}+y_{i}\cdot\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},1})$
$=4\sigma^{-2}kd^{-1}(-\bm{\eta}^{\top}\bar{\xi}_{{\mathcal{T}},1}+\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},1})$
@@
Finally, let us calculate ${\mathbf{w}}_{\mathrm{\text{DD}}}^{\top}(\bm{\mu}-\bm{\eta})$,
${\mathbf{w}}_{\mathrm{\text{DD}}}^{\top}(\bm{\mu}-\bm{\eta})$ $=2(r_{\mu}^{2}-\sigma r_{\mu}t)^{-1}r^{2}_{\mu}+4\sigma^{-2}kd^{-1}\bm{\eta}^{\top}\bar{\xi}_{{\mathcal{T}},1}$
@@
${\mathbf{w}}_{\mathrm{\text{DD}}}^{\top}(\bm{\mu}-\bm{\eta})>0$
@@
$\|{\mathbf{w}}_{\mathrm{\text{DD}}}\|=\left\|\frac{2}{r^{2}_{\mu}-\sigma r_{\mu}t}\cdot\bm{\mu}-\frac{4k}{\sigma^{2}d}\bar{\xi}_{{\mathcal{T}},1}\right\|.$
@@
$\|{\mathbf{w}}_{\mathrm{\text{DD}}}\|\leq\frac{2}{r^{2}_{\mu}-\sigma r_{\mu}t_{1}}\cdot\|\bm{\mu}\|+\frac{4k}{\sigma^{2}d}\cdot\|\bar{\xi}_{{\mathcal{T}},1}\|.$
We know that $\|\bm{\mu}\|=r_{\mu}$ and $\bar{\xi}_{{\mathcal{T}},1}\sim\mathcal{N}(0,\frac{\sigma^{2}}{k}I_{d})$. Hence, using the concentration bound in 22, we have with probability at least $1-\exp{\left(-\frac{t^{2}}{4}\right)}$ that
@@
$\leq\frac{2}{r_{\mu}-\sigma t}+8\sigma^{-1}\sqrt{\frac{k}{d}}\text{ \hskip 11.38109pt for $t\leq\frac{3}{2}\sqrt{d}$}$
@@
$\|{\mathbf{w}}^{*}\|\geq 2\|\bm{\eta}(\frac{2|U|-k}{|U|})+\bar{\xi}_{U}\|^{-1}$
$\|{\mathbf{w}}^{*}\|\geq 2(\|\bm{\eta}\|(\frac{2|U|-k}{|U|})+\|\bar{\xi}_{U}\|)^{-1}$
Similarly, using concentration bound in 22 for $\|\bar{\xi}_{U}\|$, we get with probability at least $1-\exp{\left(-\frac{t^{2}}{4}\right)}$,
@@
$C_{d}\log(\frac{c_{d}}{\delta})$ $\leq\frac{\sqrt{d}}{4\sqrt{N_{{\mathcal{S}}}+N_{{\mathcal{T}}}}},$
$\frac{1}{2\sqrt{N_{{\mathcal{T}}}}}$ $\leq r_{\mu}\leq\min\left(2\sqrt{N_{{\mathcal{T}}}}(1-5\tau),16\sqrt{N_{{\mathcal{T}},1}}\tau\right),$
$r_{\eta}$ $\geq\frac{1}{2\sqrt{N_{{\mathcal{T}},1}}},$
$\frac{3}{r_{\eta}}+\frac{3-5\tau}{r_{\mu}}$ $\leq\frac{3}{8}\sqrt{N_{{\mathcal{T}},1}}-2$
then with probability at least $1-\delta$ over the drawing of $(S_{\mathcal{S}},S_{\mathcal{T}})$ as described in problem, $AU-ROC(\hat{{\mathbf{w}}}_{\mathrm{\text{color}}})\geq 0.9$ as defined in first part of eq.˜2 is at least $Q\left(\frac{-\tau\sqrt{d}}{\sqrt{2}\|\hat{{\mathbf{w}}}_{\mathrm{\text{color}}}\|}\right)\geq Q\left(\frac{-\tau\sqrt{d}}{\sqrt{2}(\frac{3}{r_{\eta}}+\frac{3-5\tau}{r_{\mu}}+2)}\right)\geq Q\left(\frac{-\tau\sqrt{d}}{\sqrt{2}(1-\tau)}\right)$.
@@
$\mathcal{L}({\mathbf{w}},\lambda,\nu)=\|{\mathbf{w}}\|^{2}-\lambda_{{\mathcal{S}}}^{\top}(\bm{X}_{{\mathcal{S}}}{\mathbf{w}})+\lambda_{{\mathcal{T}}}^{\top}(\mathbf{1}_{{\mathcal{T}}}+\bm{X}_{{\mathcal{T}}}{\mathbf{w}})-\nu\cdot({\mathbf{w}}^{\top}(\bm{\mu}-\bm{\eta})+\tau)$
@@
$\nabla_{{\mathbf{w}}}\mathcal{L}({\mathbf{w}},\lambda,\nu)=2{\mathbf{w}}-\bm{X}^{\top}_{{\mathcal{S}}}\lambda_{{\mathcal{S}}}+\bm{X}^{\top}_{{\mathcal{T}}}\lambda_{{\mathcal{T}}}-\nu\cdot(\bm{\mu}-\bm{\eta})=0$
$\Leftrightarrow{\mathbf{w}}=\frac{1}{2}\left(\nu\cdot(\bm{\mu}-\bm{\eta})+\bm{X}^{\top}_{{\mathcal{S}}}\lambda_{{\mathcal{S}}}-\bm{X}^{\top}_{{\mathcal{T}}}\lambda_{{\mathcal{T}}}\right)$
@@
$\mathcal{L}(\lambda,\nu)=-\nu\cdot\tau+\lambda_{{\mathcal{T}}}^{\top}\mathbf{1}_{{\mathcal{T}}}-\frac{1}{4}\|\nu\cdot(\bm{\mu}-\bm{\eta})+\bm{Z}^{\top}\lambda\|^{2}$
Consider setting $\lambda=\frac{\nu}{|U|}\mathbf{1}_{S_{{\mathcal{T}}}}$, where $U$ has randomly drawn $|U|$ ($|U|=N_{\mathcal{T}}$) examples from $S_{{\mathcal{T}}}$. Also set $\nu=\frac{(1-\tau)}{\|\bar{\xi}_{S_{{\mathcal{T}}}}\|^{2}}$. We have
$\bm{Z}^{\top}\lambda$ $=-\nu\cdot\bm{\mu}+\nu\bar{\xi}_{S_{{\mathcal{T}},0}}+\nu\cdot\bm{\eta}+\nu\bar{\xi}_{S_{{\mathcal{T}},1}}$
$=\nu(\bm{\eta}-\bm{\mu})+\nu\bar{\xi}_{S_{{\mathcal{T}}}}$
@@
${\mathbf{w}}_{\mathrm{color}}=\alpha\bm{\eta}+(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}})\bm{\mu}-\beta\bar{\xi}_{S_{{\mathcal{T}},0}}$
${\mathbf{w}}_{\mathrm{color}}=2\frac{\bm{\eta}}{r^{2}_{\eta}}+(\frac{2}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}})\bm{\mu}-\frac{1}{\sigma^{2}}\cdot\frac{|U|-k}{d}\bar{\xi}_{S_{{\mathcal{T}},0}}$
${\mathbf{w}}_{\mathrm{color}}=2\frac{\bm{\eta}}{r^{2}_{\eta}}+(\frac{2}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}})\bm{\mu}-(|U|-k)\bar{\xi}_{S_{{\mathcal{T}},0}}$
@@
${\mathbf{w}}_{\mathrm{color}}^{\top}(\bm{\eta}-\xi_{i})=\alpha r^{2}_{\eta}+\beta\xi_{i}^{\top}\bar{\xi}_{S_{{\mathcal{T}},0}}>1$
@@
${\mathbf{w}}_{\mathrm{color}}^{\top}(-y_{i}\bm{\eta}+y_{i}\xi_{i})$ $={\mathbf{w}}_{\mathrm{color}}^{\top}(\bm{\eta}-\xi_{i})$
$=\alpha r^{2}_{\eta}-\alpha\bm{\eta}^{\top}\xi_{i}+\beta\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},0}$
We know that $|\bm{\eta}^{\top}\xi_{i}|<\sigma tr_{\eta}$ w.p. at least $1-2\exp\left(-ct^{2}/2\right)$ and $|\xi_{i}\bar{\xi}_{{\mathcal{T}},0}|\leq\sigma^{2}t\sqrt{\frac{d}{|U|-k}}$ w.p. at least $1-2\exp(-ct^{2})$ for $t\leq d\sqrt{|U|}\sigma^{2}$. Hence, w.p. at least $1-2\exp\left(-ct^{2}/2\right)-2\exp(-ct^{2})$, we can say that,
${\mathbf{w}}_{\mathrm{color}}^{\top}(\bm{\eta}-\xi_{i})\geq\alpha r^{2}_{\eta}-\alpha\sigma tr_{\eta}-\beta\sigma^{2}t\sqrt{\frac{d}{|U|-k}}$
@@
${\mathbf{w}}_{\mathrm{color}}^{\top}(\bm{\mu}+\xi_{i})=(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}})r^{2}_{\mu}+\beta\xi_{i}^{\top}\bar{\xi}_{S_{{\mathcal{T}},0}}>1$
@@
${\mathbf{w}}_{\mathrm{color}}^{\top}(-y_{i}\bm{\mu}+y_{i}\xi_{i})$ $={\mathbf{w}}_{\mathrm{color}}^{\top}(\bm{\mu}-\xi_{i})$
$=\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)(\bm{\mu}^{\top}\bm{\mu}-\bm{\mu}^{\top}\xi_{i})-\beta(\bm{\mu}^{\top}\bar{\xi}_{{\mathcal{T}},0}-\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},0})$
We know that $|\bm{\mu}^{\top}\xi_{i}|\leq\sigma tr_{\mu}$ w.p. at least $1-2\exp\left(-ct^{2}/2\right)$ and $|\xi_{i}^{\top}\bar{\xi}_{{\mathcal{T}},0}|\geq\frac{\sigma^{2}d}{2(|U|-k)}$ w.p. at least $1-2\exp(-ct^{2})$ for $t\leq\sigma^{2}$ and $t\leq\frac{\sqrt{d}}{2(2+\sqrt{|U|-1})}$. We also know that $|\bm{\mu}^{\top}\bar{\xi}_{{\mathcal{T}},0}|\leq\frac{\sigma tr_{\mu}}{|U|-k}$ w.p. at least $1-2\exp(-ct^{2}/2)$ Hence, w.p. at least $1-4\exp\left(-ct^{2}/2\right)-2\exp(-ct^{2})$, we can say that,
${\mathbf{w}}_{\mathrm{color}}^{\top}(-y_{i}\bm{\mu}+y_{i}\xi_{i})$ $\geq\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)(r^{2}_{\mu}-\sigma tr_{\mu})+\beta\left(-\frac{\sigma tr_{\mu}}{|U|-k}+\frac{\sigma^{2}d}{2(|U|-k)}\right)$
$\geq\frac{r^{2}_{\mu}}{2}\left(\frac{\alpha r^{2}_{\eta}}{r^{2}_{\mu}}-\frac{5\tau}{r^{2}_{\mu}}\right)+\beta\left(\frac{\sigma^{2}d-r^{2}_{\mu}}{2(|U|-k)}\right)\text{ \hskip 11.38109pt for $r_{\mu}\geq 2\sigma t$}$